# Code Execution as Grounded Supervision for LLM Reasoning

### v1

@@ 2.2 Execution Trace Acquisition
We employ a Python debugging tool called Snoop (Hall, 2024) as the execution tracing tool, which records detailed line-by-line execution signals, including function calls and returns, executed lines of code, and the updated local variable values. An example of a Snoop-generated execution trace is provided in Figure 1.222A more illustrative example of an execution trace is provided in Appendix B The resulting dataset is denoted as $D_{\text{trace}}=(\mathbf{x},\mathbf{s}_{\text{trace}},\mathbf{y})$, which contains verifiably accurate execution trace $\mathbf{s}_{\text{trace}}$ and the correct final output $\mathbf{y}$ for each instance, grounded in a verifiable code executor.
@@ 3.2 Experiment Results
The experiment results of Qwen3-4B is presented in Table 1.444The experiment result of Qwen3-8B is presented in Table 6 Overall, our method outperforms all baselines, demonstrating its effectiveness. The Code Generation and Raw Execution Trace baselines perform poorly across the board. Although both are derived from code-based supervision, they suffer from a misalignment with natural language reasoning formats. In particular, training on raw traces or code generation alone does not equip the model with the ability to produce step-by-step reasoning in natural language, leading to limited transfer and degraded performance—especially in reasoning-heavy tasks.
@@ 3.3 Ablation Studies
Table 2 shows the accuracy of both the final outputs and intermediate reasoning steps for our method and CodeI/O. Our method achieves higher accuracy than CodeI/O on both metrics, with a particularly larger margin in intermediate step accuracy. This demonstrates that our method produces more accurate reasoning steps, leading to the correct final output, as it is grounded in reliable code execution.
Reducing Repetition and Overthinking. To evaluate the generation efficiency of the models trained on our method, we compare token lengths in Table 3. We show that models trained with our approach generate approximately 20% fewer tokens for Qwen3-4B and 30% fewer tokens for Qwen3-8B, compared to No Training baseline. [^5] We include Qwen3-8B results in Table 5 Moreover, our method substantially reduces instances where the model reaches the maximum token limit due to overthinking or repetitive output. Importantly, this reduction in token length does not compromise performance, as demonstrated in Section 3.2.
@@ 3.4 Case Study
To further examine our data, we present two examples of reasoning trace in Appendix D and Appendix D. The first example demonstrates a case where the CodeI/O solution produces the correct final output despite an error in an intermediate step—specifically, it incorrectly lists the permutation of "hrf" in step 2—while our solution correctly completes all intermediate steps. The second example shows a case where both the intermediate reasoning and the final output are incorrect in the CodeI/O solution, due to an incorrect formula used at the beginning. In contrast, our solution produces correct calculations throughout, guided by the execution trace.
@@ A.1 Training
We train all models using LLaMA-Factory Zheng et al. (2024b) on 8 NVIDIA A100-SXM4-40G GPUs. We use full parameter fine-tuning across all the models in our experiment. Training hyperparameters are detailed in Table 4.
@@ A.5 Prompt Templates
We present the prompt templates used in the experiment in Table 8.
@@ Appendix B Execution Trace Example
Appendix D presents an example of executable code alongside its execution trace generated by Snoop. To enable tracing, the @snoop decorator must be applied to the main entry function. The resulting execution trace includes function calls and return values, executed lines annotated with line numbers, and updated variable values.
@@ Appendix C Additional Experiment Results
We present the experiment results for Qwen3-8B in Table 6. Similar to the results of Qwen3-4B, our method overall outperforms all the baselines with particular strength in the coding benchmark.
Additionally, we present token length analysis of Qwen3-8B on LiveBench in Table 5
@@ Appendix D Licenses
import snoop # Import Snoop library @snoop # Add the decorator to trace the function def main_solution(num): if num < 0: return ’-’ + str(main_solution(-num)) elif num < 7: return str(num) else: return str(main_solution(num // 7)) + str(num % 7) main_solution(num=100) # Function call $>$ Call to main_solution # Function call ...... num = 100 # Input 38 | def main_solution(num): 39 | if num < 0: # Executed line of code 41 | elif num < 7: 44 | return str(main_solution(num // 7)) + str(num % 7) $>$ Call to main_solution ...... num = 14 # Variable value 38 | def main_solution(num): 39 | if num < 0: 41 | elif num < 7: 44 | return str(main_solution(num // 7)) + str(num % 7) $>$ Call to main_solution ...... num = 2 38 | def main_solution(num): 39 | if num < 0: 41 | elif num < 7: 42 | return str(num) $<$ Return value from main_solution: ’2’ 44 | return str(main_solution(num // 7)) + str(num % 7) $<$ Return value from main_solution: ’20’ 44 | return str(main_solution(num // 7)) + str(num % 7) $<$ Return value from main_solution: ’202’ # Return

### v2

@@ 2.2 Execution Trace Acquisition
We employ a Python debugging tool called Snoop (Hall, 2024) as the execution tracing tool, which records detailed line-by-line execution signals, including function calls and returns, executed lines of code, and the updated local variable values. An example of a Snoop-generated execution trace is provided in Figure˜1.222A more illustrative example of an execution trace is provided in Appendix B The resulting dataset is denoted as $D_{\text{trace}}=(\mathbf{x},\mathbf{s}_{\text{trace}},\mathbf{y})$, which contains verifiably accurate execution trace $\mathbf{s}_{\text{trace}}$ and the correct final output $\mathbf{y}$ for each instance, grounded in a verifiable code executor.
@@ 3.2 Experiment Results
The experiment results of Qwen3-4B is presented in Table˜1.444The experiment result of Qwen3-8B is presented in Table 6 Overall, our method outperforms all baselines, demonstrating its effectiveness. The Code Generation and Raw Execution Trace baselines perform poorly across the board. Although both are derived from code-based supervision, they suffer from a misalignment with natural language reasoning formats. In particular, training on raw traces or code generation alone does not equip the model with the ability to produce step-by-step reasoning in natural language, leading to limited transfer and degraded performance—especially in reasoning-heavy tasks.
@@ 3.3 Ablation Studies
Table˜2 shows the accuracy of both the final outputs and intermediate reasoning steps for our method and CodeI/O. Our method achieves higher accuracy than CodeI/O on both metrics, with a particularly larger margin in intermediate step accuracy. This demonstrates that our method produces more accurate reasoning steps, leading to the correct final output, as it is grounded in reliable code execution.
Reducing Repetition and Overthinking. To evaluate the generation efficiency of the models trained on our method, we compare token lengths in Table˜3. We show that models trained with our approach generate approximately 20% fewer tokens for Qwen3-4B and 30% fewer tokens for Qwen3-8B, compared to No Training baseline. [^5] We include Qwen3-8B results in Table 5 Moreover, our method substantially reduces instances where the model reaches the maximum token limit due to overthinking or repetitive output. Importantly, this reduction in token length does not compromise performance, as demonstrated in Section˜3.2.
@@ 3.4 Case Study
To further examine our data, we present two examples of reasoning trace in Appendix˜D and Appendix˜D. The first example demonstrates a case where the CodeI/O solution produces the correct final output despite an error in an intermediate step—specifically, it incorrectly lists the permutation of "hrf" in step 2—while our solution correctly completes all intermediate steps. The second example shows a case where both the intermediate reasoning and the final output are incorrect in the CodeI/O solution, due to an incorrect formula used at the beginning. In contrast, our solution produces correct calculations throughout, guided by the execution trace.
@@ 5 Conclusion
Acknowledgment
We appreciate the reviewers for their insightful comments and suggestions. This work was partly supported by the Amazon Nova Trusted AI Prize, the NSF of the United States Grants ITE 2333736 and OAC 2531126, and the DARPA FoundSci Grant HR00112490370.
@@ A.1 Training
We train all models using LLaMA-Factory Zheng et al. (2024b) on 8 NVIDIA A100-SXM4-40G GPUs. We use full parameter fine-tuning across all the models in our experiment. Training hyperparameters are detailed in Table˜4.
@@ A.5 Prompt Templates
We present the prompt templates used in the experiment in Table˜8.
@@ Appendix B Execution Trace Example
Appendix˜D presents an example of executable code alongside its execution trace generated by Snoop. To enable tracing, the @snoop decorator must be applied to the main entry function. The resulting execution trace includes function calls and return values, executed lines annotated with line numbers, and updated variable values.
@@ Appendix C Additional Experiment Results
We present the experiment results for Qwen3-8B in Table˜6. Similar to the results of Qwen3-4B, our method overall outperforms all the baselines with particular strength in the coding benchmark.
Additionally, we present token length analysis of Qwen3-8B on LiveBench in Table˜5
@@ Appendix D Licenses
import snoop # Import Snoop library @snoop # Add the decorator to trace the function def main_solution(num): if num < 0: return ’-’ + str(main_solution(-num)) elif num < 7: return str(num) else: return str(main_solution(num // 7)) + str(num % 7) main_solution(num=100) # Function call >>> Call to main_solution # Function call ...... num = 100 # Input 38 | def main_solution(num): 39 | if num < 0: # Executed line of code 41 | elif num < 7: 44 | return str(main_solution(num // 7)) + str(num % 7) >>> Call to main_solution ...... num = 14 # Variable value 38 | def main_solution(num): 39 | if num < 0: 41 | elif num < 7: 44 | return str(main_solution(num // 7)) + str(num % 7) >>> Call to main_solution ...... num = 2 38 | def main_solution(num): 39 | if num < 0: 41 | elif num < 7: 42 | return str(num) <<< Return value from main_solution: ’2’ 44 | return str(main_solution(num // 7)) + str(num % 7) <<< Return value from main_solution: ’20’ 44 | return str(main_solution(num // 7)) + str(num % 7) <<< Return value from main_solution: ’202’ # Return